OpenStudy (anonymous):

Solve the system of equations graphically. If the system does not have a single ordered pair as a solution, state whether the system is inconsistent of dependent. 9x - 2y = 30 5x + y = 24

OpenStudy (anonymous):

Okie dokie I want to say its (3, 4) correct me if wrong :/

OpenStudy (anonymous):

Just by plugging those numbers into the original equations, your answer is wrong. Try again.

OpenStudy (anonymous):

Multiply the second equation by 2, and then add the system. See what happens?

OpenStudy (anonymous):

okay :/

OpenStudy (anonymous):

Do it up @gezimbasha

OpenStudy (anonymous):

(4, -3) :s ?

OpenStudy (anonymous):

hold up the question ask 9x - 2y = 30 5x + y = 24 its supposed to look like this 9x - 2y = 30 5x + 4/3 y = 24

OpenStudy (anonymous):

Ooh, that changes things alot. It makes your first answer correct. Great job!!

OpenStudy (anonymous):

Lol thank you I was like there is something wrong here. x) ha dumby me thanks

OpenStudy (anonymous):

Yeah me too, it gave answers in decimal which isn't usually what teachers do. They're bad, but not that bad ;)